1.How many millimoles of solute are contained in
a.2.95ML of 0.0789M KH2PO4?
b.0.2011L of 0.0564M HgCl2?

2.Calculate the PH of a solution that contains 2.0% (w/w)NaOH and has a density of 1.022g/ML.
1
Expert's answer
2021-08-02T12:26:32-0400

Solution:

(1):

1 mmol = 0.001 mol

Moles = Molarity × Volume

Therefore,

a)

Moles of KH2PO4 = (0.0789 M) × (2.95 mL) × (1 L / 1000 mL) = 0.000233 mol

Millimoles of KH2PO4 = (0.000233 mol) × (1 mmol / 0.001 mol) = 0.233 mmol

Millimoles of KH2PO4 = 0.233 mmol


b)

Moles of HgCl2 = (0.0564 M) × (0.2011 L) = 0.01134 mol

Millimoles of HgCl2 = (0.01134 mol) × (1 mmol / 0.001 mol) = 11.34 mmol

Millimoles of HgCl2 = 11.34 mmol


(2):

Determine the mass of NaOH per milliliter using the mass percentage of NaOH and the density:

(0.0200) × (1.022 g / 1 mL) = 0.02044 g NaOH mL-1

The molar mass of NaOH is 39.997 g/mol.

Therefore,

(0.02044 g NaOH / 1 mL) × (1 mol NaOH / 39.997 g NaOH) × (1000 mL / 1L) = 0.511 mol NaOH L-1

Molarity of NaOH = CM(NaOH) = 0.511 M


NaOH (sodium hydroxide) is a strong base which dissociates as follows:

NaOH → Na+ + OH-

[OH-] = CM(NaOH) = 0.511 M (according to the equation above)

We can convert between [OH-] and pOH using the following equation:

pOH = - log[OH-]

pOH = - log(0.511) = 0.29

pOH = 0.29

For any aqueous solution at 25C:

pH + pOH = 14

Thus,

pH = 14 - pOH = 14 - 0.29 = 13.71

pH = 13.71


Answers:

(1): a) 0.233 mmol KH2PO4; b) 11.34 mmol HgCl2

(2): pH = 13.71